The formula for calculating the amount of money returned for an initial deposit into a bank account or CD (certificate of deposit) is given by

A is the amount of the return.
P is the principal amount initially deposited.
r is the annual interest rate (expressed as a decimal).
n is the number of compound periods in one year.
t is the number of years.
If a bank compounds continuously, then the formula used is
where e is a constant and equals approximately 2.7183.
Calculate A with continuous compounding. Round your answer to the hundredth's place.
Answer:
Show work in this space:

e) A commonly asked question is, “How long will it take to double my money?” At 8% interest rate and continuous compounding, what is the answer? Round your answer to the hundredth's place.
Answer:
Show work in this space:

Answer by stanbon(75887)

“How long will it take to double my money?” At 8% interest rate and continuous compounding
A(t) = Ao*e^(rt)
Ao is the amount invested
--------------
Your Problem:
2Ao = Aoe^rt
2 = e^rt
Take the natural log of both sides to get:
ln2 = rt
ln2 = 0.08t
t = ln2 / 0.08
t = 8.6643 years
-------------
Cheers,
Stan H.